OpenStudy (anonymous):

graph the linear equation by solving for y first x+5y=20

OpenStudy (anonymous):

get y by itself

OpenStudy (anonymous):

Okay to solve for y x+5y=20 subtract x on both sides -x -x ____________ 5y= -x+20 divide 5 on both sides 5y/5= (-x+20)/ 5 y= -x/5+4

OpenStudy (anonymous):

4 is your y intercept so you will plot 4 on the graph

OpenStudy (anonymous):

Next you have -x/5 All the this (-x/5) says is that you go down 1 and over 5, do you understand??

OpenStudy (anonymous):



OpenStudy (anonymous):

That was correct (and nicely explained!) until the last step. The graph of y = -x/5 + 4 doesn't intersect the x-axis at (5,0)... when y = 0, the equation is 0 = -x/5 + 4 so x/5 = 4 making x = 20 for y = 0. The graph goes down 1 unit for every 5 it moves to the right. So to move down from the y-intercept of (0,4) to the x-axis, it has to drop 4 units down... so it moves 4(5) = 20 units to the right.

OpenStudy (anonymous):